# ThermoNeRF: Multimodal Neural Radiance Fields for Thermal Novel View Synthesis

### v1

@@ Abstract
Thermal scene reconstruction exhibit great potential for applications across a broad spectrum of fields, including building energy consumption analysis and non-destructive testing. However, existing methods typically require dense scene measurements and often rely on RGB images for 3D geometry reconstruction, with thermal information being projected post-reconstruction. This two-step strategy, adopted due to the lack of texture in thermal images, can lead to disparities between the geometry and temperatures of the reconstructed objects and those of the actual scene. To address this challenge, we propose ThermoNeRF, a novel multimodal approach based on Neural Radiance Fields, capable of rendering new RGB and thermal views of a scene jointly. To overcome the lack of texture in thermal images, we use paired RGB and thermal images to learn scene density, while distinct networks estimate color and temperature information. Furthermore, we introduce ThermoScenes, a new dataset to palliate the lack of available RGB+thermal datasets for scene reconstruction. Experimental results validate that ThermoNeRF achieves accurate thermal image synthesis, with an average mean absolute error of C, an improvement of over 50% compared to using concatenated RGB+thermal data with Nerfacto, a state-of-the-art NeRF method. Code is available at https://github.com/SchindlerEPFL/thermo-nerf.
@@
Thermal Imaging Neural Radiance Fields 3D Reconstruction Multimodality1 Introduction
Since their introduction by [40], Neural Radiance Fields (NeRFs) have achieved great success in 3D reconstruction and novel view synthesis and have been extended for diverse tasks such as depth estimation [9, 51], semantic segmentation [62, 34] and object detection [56, 24]. While most NeRF models learn an implicit representation of a scene from a sparse set of RGB images, NeRFs have also been successfully extended to other sensor modalities such as depth views [63], near-infrared and multi-spectral images [44], LiDAR point clouds [59, 63] and even audio signals [10]. Moreover, it has been demonstrated that NeRFs can learn from more than a single modality at a time—for example, [44] combine near-infrared and RGB images to learn a more accurate representation of the scene than when using only one modality. However, multimodal NeRFs also often depend on the presence of salient or similar features between the modalities [44].
In this paper, we focus on multimodal NeRF using RGB and thermal imaging. Thermal cameras capture the radiation emitted by objects as a function of their temperature. At ambient temperature ranges, objects radiate in the mid- and far-infrared region of the electromagnetic spectrum. Since thermal radiation is scattered and reflected from various points within a scene, thermal images naturally lack texture and appear soft—i.e., with a low edge contrast—an effect known as ghosting [1]. Hence, despite the benefits of using 3D thermal representations in infrastructure inspection and monitoring [28, 22], non-destructive testing [22], and agriculture [43], such models are primarily constructed using photogrammetry, which requires large amounts of RGB data or advanced equipment [39, 15, 37]. Moreover, thermal images are often not considered during the geometric reconstruction, and temperature information is projected post-reconstruction [45], leading to potential discrepancies between the model and actual temperatures.
Additionally, the development of thermal scene reconstruction algorithms is hindered by the lack of benchmark datasets specifically designed for this task. Instead, the majority of existing datasets that include thermal images are focused on dynamic and changing environments for a variety of perception tasks, such as object detection [26, 18, 19], object tracking [5], face or object recognition [29, 12], semantic segmentation [52], motion estimation [14], SLAM [57] and autonomous driving [13, 32].
@@
To address the gaps outlined above, we propose ThermoNeRF (Thermographic NeRF), the first multimodal NeRF model capable of rendering unseen views in both RGB and thermal modalities jointly (see Fig. 1). The design of our approach is guided by the specific properties of thermal images. ThermoNeRF uses a shared density MLP that leverages the visual features of the RGB modality to learn the geometry of the scene while ensuring consistency with the thermal measurements. Color and thermal information are learned through decoupled MLPs, to prevent the influence of colors on the estimated temperatures, and vice-versa. In addition, we also introduce ThermoScenes, a new paired thermal-RGB dataset comprising ten diverse scenes—six indoor and four outdoor scenes— for 3D scene reconstruction and novel view synthesis.
@@
We propose ThermoNeRF, the first multimodal NeRF capable of rendering both thermal and RGB views jointly.
We conduct extensive experiments to validate how ThermoNeRF’s architecture allows the RGB modality to guide the density estimation for thermal reconstruction while preventing interactions between color and estimated temperatures.
We provide ThermoScenes, a new RGB+thermal image dataset for 3D scene reconstruction and novel view synthesis, featuring ten scenes with diverse temperature ranges, object types and sizes.
Finally, we present a comprehensive evaluation of our method’s performance, covering both temperature estimation and reconstruction quality on unseen poses. Our results demonstrate improved temperature estimation with no loss in reconstruction fidelity when compared to models trained using only thermal images or concatenated RBG+thermal images as input.
@@ 2.1 Thermal Computer Vision
Thermal cameras measure thermal radiation emitted by any matter with a temperature above absolute zero Kelvin. While most practical temperature ranges fall within the infrared (IR) region of the electromagnetic spectrum [25], thermal cameras operate in the mid- to far-IR spectrum, unlike near-infrared (NIR) sensors, which work closer to the visible range. It is important to note that thermal cameras also detect thermal radiations that are scattered and reflected by objects and the environment, resulting in thermal images that are naturally textureless and soft—a phenomenon known as the ghosting effect [1]. While researchers have explored object detection [17] and tracking [6] using infrared sensors, as well as the analysis of thermal images [27, 36, 35], due to their soft and textureless nature, working with thermal images is a challenge [1].
Hence, to model scenes with thermal data, early studies predominantly relied on range measurements—e.g. from LiDAR [31], depth cameras [53, 48] or even a combination of both sensors [8]—to construct 3D models, onto which thermal information was subsequently projected. However, those approaches require a time-consuming and error-prone manual calibration of sensor transformations, their precision depends on the accuracy of mapping algorithms, and in the case of depth cameras, are limited by their range. To remove the dependence on range measurements, [39, 15] use photogrammetry to construct 3D models from RGB images. While these methods achieve high model accuracy, they depend on a dense set of RGB and thermal images.
While previous works presented in this section rely on range or RGB measurements to build the 3D geometry of a scene, [11, 49] used thermal images for 3D reconstruction and temperature estimation. However, their results are limited to thermal images with high contrast and do not generalize to large structures.
@@ 2.2 NeRF and Multimodality
Traditional methods in 3D reconstruction, such as Structure from Motion [46] or Multi-View Stereo [47], depend on dense input data and scenes rich in geometric features. The seminal work by [40] introduced NeRF, enabling the learning of implicit representations of 3D scenes from sparse sets of RGB images for novel view synthesis. Subsequent research efforts have extended and refined the NeRF framework, addressing challenges related to scalability [38], level of detail [2, 4] and efficiency [58].
@@
Closely related to our work on thermal imaging, [44] integrates near-infrared and RGB images to create a composite NIR-RGB 3D scene. However, NIR images share similar features with RGB images, as they are closer to the visible spectrum compared to thermal images.
To the best of our knowledge, ThermoNeRF is the first approach to leverage a NeRF framework for thermal scene representation and novel view synthesis from a sparse set of images. Unlike existing methods that rely on multimodal data with salient visual features, ThermoNeRF combines RGB and thermal features to learn consistent scene geometries across both sensor modalities, while avoiding contamination from each modality in the rendered views.
@@ 4 ThermoNeRF
In this section, we present ThermoNeRF, a NeRF model capable of learning an implicit scene representation in both thermal and RGB, from a sparse set of RGB and thermal images. To synthesize realistic and accurate RGB and thermal novel views, ThermoNeRF retains geometric features captured by the RGB modality, while ensuring that temperature estimates are independent of scene color variations. Our design choices are guided by two properties of thermal images: 1) Thermal images are inherently soft and textureless. 2) Temperatures are independent of the viewing direction.
Refer to Fig. 2 for a flowchart of the method; implementation of the model and evaluations can be found online111https://github.com/SchindlerEPFL/thermo-nerf.
@@ 4.1 Thermal Image Rendering
As depicted in Fig. 3, while RGB information can be view-dependent due to non-Lambertian effects, thermal cameras measure thermal radiation and convert it to temperature, which is invariant with respect to the viewing direction. Therefore, ThermoNeRF uses separate MLPs for each modality: predicts temperatures along each ray, while predicts RGB values. Since temperatures are independent of the viewing direction, only takes the intermediate features of as input, while receives , and —as presented in Sec. 3.
Furthermore, as seen in the thermal images in Fig. 3, the ghosting effect makes thermal images soft and textureless. Consequently, although textures and sharp geometric features in RGB images allow NeRF models to accurately estimate scene density, NeRF struggles to learn densities solely from thermal images—see Fig. 3(b) for experimental results of training Nerfacto on thermal images only. Therefore, in ThermoNeRF, a single density is shared by the modalities, enabling the network to construct a geometric representation informed by RGB information while ensuring geometric consistency with thermal modality information.
@@
where t denotes the predicted temperature values.
@@ 4.2 Loss Functions
We use separate reconstruction loss functions for each modality. The final reconstruction loss is expressed as the sum of the Mean Square Error (MSE) losses of the RGB and thermal outputs, respectively denoted and . Similar to Nerfacto, we add the interlevel and distortion losses, which were initially used in MIP-NeRF 360 [3], to optimize the proposal sampler and to reduce distortions respectively. Therefore, our final loss is defined as follows:
@@ 5 ThermoScenes Dataset
To evaluate RGB+thermal reconstruction and novel view synthesis methods, we have collected ThermoScenes, a new dataset comprising paired RGB and thermal images for ten scenes (four outdoor and six indoor) with diverse temperature ranges. To collect this dataset, we used a FLIR One Pro LT [21], a dual RGB and thermal camera, mounted on an iPhone, capable of simultaneously capturing aligned pairs of RGB and thermal images. The FLIR One Pro LT thermal camera operates within a range of -20 °C to 120 °C, with a thermal accuracy of 3 °C. In our study, evaluating the quality of 3D thermal reconstruction requires not only understanding the thermal accuracy but also the thermal precision of the camera. We assess the precision to be approximately C. To calculate the precision of our thermal camera, we fix the camera at a position and capture 20 consecutive images of a RaspberryPi connected to power for at least 10 minutes to ensure thermal stability. Our precision is then the average standard deviation of the temperatures of each pixel across the 20 images.
Furthermore, the FLIR software generates thermal images by overlaying thermal information onto RGB features—such as edges—using Multi-Spectral Dynamic Imaging (MSX) [55]. For this dataset, we extract raw thermal data from MSX images using the FLIR Image Extractor library [20], and disregard the MSX-enhanced images.
@@
*In one scene, the sensor outputs values outside of its operating range, which are detected on the clear sky. Due to the variations of atmospheric radiation in the sky, the readings of thermal cameras are subject to systematic inaccuracies [30]. We clip these values to the minimum of the operating range of the camera (-20°C). However, our data can be post-processed for more accurate temperature estimates in the sky [30, 61].
Tab. 1 provides a summary of the collected ThermoScenes dataset, where each scene is illustrated by two example views in RGB and thermal respectively. The test set for each scene comprises one-eighth of the total number of images, with test views uniformly sampled across the collected camera poses. The minimum and maximum measured temperatures are also reported. More images for all ten scenes, along with a thorough explanation of each scene, are included in the Dataset Description section of the supplementary material. ThermoScenes is publicly available222https://drive.google.com/drive/folders/1Je6pyQUV3keCStn1yLsyyPEdlccm70B3.

### v2

# ThermoNeRF: Joint RGB and Thermal Novel View Synthesis for Building Facades using Multimodal Neural Radiance Fields

@@ Abstract
Thermal scene reconstruction holds great potential for various applications, such as analyzing building energy consumption and performing non-destructive infrastructure testing. However, existing methods typically require dense scene measurements and often rely on RGB images for 3D geometry reconstruction, projecting thermal information post-reconstruction. This can lead to inconsistencies between the reconstructed geometry and temperature data and their actual values. To address this challenge, we propose ThermoNeRF, a novel multimodal approach based on Neural Radiance Fields that jointly renders new RGB and thermal views of a scene, and ThermoScenes, a dataset of paired RGB+thermal images comprising 8 scenes of building facades and 8 scenes of everyday objects. To address the lack of texture in thermal images, ThermoNeRF uses paired RGB and thermal images to learn scene density, while separate networks estimate color and temperature data. Unlike comparable studies, our focus is on temperature reconstruction and experimental results demonstrate that ThermoNeRF achieves an average mean absolute error of C and C for temperature estimation in buildings and other scenes, respectively, representing an improvement of over 50% compared to using concatenated RGB+thermal data as input to a standard NeRF. Code and dataset are available online.
@@
[…]
We propose ThermoNeRF, a method to synthesize novel RGB+thermal views.
We introduce ThermoScenes, a novel dataset with 16 scenes (with 8 building facades).
We show that separating RGB and thermal fields prevents information leakage.
Joint learning of the density improves density estimations for both modalities.
1 Introduction
Thermal scene reconstruction offers significant potential for a wide range of applications, including building energy consumption analysis [1], non-destructive testing [2], infrastructure inspection and monitoring [2, 3], and optimization of the cost-effectiveness of building facade retrofit [4]—notably, building facades play a crucial role in regulating the energy flowing in and out of buildings [5]. However, while thermal cameras capture radiation emitted by objects as a function of temperature, thermal images inherently lack texture and exhibit low edge contrast due to scattering and reflection from various surfaces, resulting in an effect known as ghosting [6]. Consequently, thermal images are often not considered during geometric reconstruction, with 3D thermal models predominantly constructed using photogrammetry with large amounts of RGB images [7, 8, 9]. In these models, temperature information is merely projected onto the reconstructed geometry post-reconstruction [10], leading to discrepancies between the geometry and the temperature data. The development of thermal scene reconstruction algorithms also faces challenges due to the lack of benchmark datasets specifically designed for this task, especially in the context of building facade reconstruction.
Recently, Neural Radiance Fields (NeRFs) [11] have achieved great success in 3D reconstruction and novel view synthesis by learning an implicit representation of a scene from a set of sparse RGB images. While most NeRF models focus on learning from RGB images, they have also been extended to other sensor modalities such as hyperspectral images [12], acoustic signals [13], point clouds [14, 15], or thermal images [16, 17]. However, prior research on NeRF utilizing thermal images has either relied on high-resolution thermal images captured with costly equipment or emphasized the visual quality of the reconstruction in either low-light conditions [18] or against occlusions [16], without prioritizing the precision of the estimated temperature which hinders their practical use in different applications.
@@ 1 Introduction
To bridge these research gaps, we propose (1) ThermoNeRF (Thermographic NeRF), a multimodal NeRF model capable of rendering unseen views in both RGB and thermal modalities with a focus on temperature rendering quality and not only image quality (see Fig. 1) and (2) ThermoScenes, the first paired RGB+thermal dataset comprising sixteen diverse scenes—eight building facade scenes and eight everyday objects scenes. The design of ThermoNeRF is guided by the specific properties of thermal images. ThermoNeRF uses a shared density MLP that leverages the visual features of the RGB modality to learn the geometry of the scene while ensuring consistency with the thermal measurements. On the other hand, color and thermal information are learned through decoupled MLPs, preventing the influence of colors on the estimated temperatures, and vice-versa. Both the dataset111https://doi.org/10.5281/zenodo.10835108 and the code222https://github.com/Schindler-EPFL-Lab/thermo-nerf (including instructions to submit new scenes to the dataset) are available online.
@@
We propose ThermoNeRF, a multimodal NeRF capable of rendering both thermal and RGB views jointly and conduct extensive experiments to validate how ThermoNeRF’s architecture allows the RGB modality to guide the density estimation for thermal reconstruction while preventing information leakage between both modalities, hence leading to accurate temperature estimations.
We provide ThermoScenes, the first RGB+thermal image dataset for 3D scene reconstruction and novel view synthesis, featuring 16 scenes with diverse temperature ranges and objects, including 8 building facades and 8 everyday object scenes.
Finally, we present a comprehensive evaluation covering both temperature estimation and reconstruction quality on unseen poses.
Our results demonstrate improved temperature estimation with no loss in reconstruction fidelity when compared to models trained using only thermal images or concatenated RBG+thermal images as input.
@@ 2.1 Thermal Computer Vision
Thermal cameras measure thermal radiation emitted by any matter with a temperature above absolute zero Kelvin. While most practical temperature ranges fall within the infrared (IR) region of the electromagnetic spectrum [19], thermal cameras operate in the mid- to far-IR spectrum, unlike near-infrared (NIR) sensors, which work closer to the visible range. Thermal cameras detect thermal radiations that are scattered and reflected by objects and the environment, resulting in thermal images that are naturally textureless and soft—a phenomenon known as the ghosting effect [6].
While researchers have explored object detection [20] and tracking [21] using infrared sensors, as well as the analysis of thermal images [22, 23, 24], due to their soft and textureless nature, working with thermal images is a challenge [6]. Hence, early studies modeling thermal data predominantly relied on range measurements—e.g. LiDAR [25], depth cameras [26, 27] or both [28]—to construct 3D models, onto which thermal information was subsequently projected. However, those approaches require a time-consuming and error-prone manual calibration of sensor transformations, their precision depends on the accuracy of mapping algorithms, and in the case of depth cameras, are limited by their range. To remove the dependence on range measurements, Maset et al. [7], De Luis-Ruiz et al. [8] use photogrammetry to construct 3D models from RGB images. While these methods achieve high model accuracy, they depend on a dense set of RGB and thermal images. On the other hand, Chen et al. [29] and Sentenac et al. [30] used thermal images for 3D reconstruction and temperature estimation but their results are limited to thermal images with high contrast and do not generalize to large structures like building facades. Hence, building thermal envelope analysis is usually done by mapping thermal data on existing models [31].
@@ 2.2 NeRF and Multimodality
Traditional methods in 3D reconstruction, such as Structure from Motion [32] or Multi-View Stereo [33], depend on dense input data and scenes rich in geometric features. The seminal work by Mildenhall et al. [11] introduced NeRF, enabling the learning of implicit representations of 3D scenes from sparse sets of RGB images for novel view synthesis. Subsequent research efforts have extended and refined the NeRF framework, addressing challenges related to scalability [34], level of detail [35, 36] and efficiency [37].
@@
Closely related to our work on thermal imaging, Poggi et al. [12] integrates near-infrared and RGB images to create a composite NIR-RGB 3D scene. However, NIR images share similar features with RGB images, as they are closer to the visible spectrum compared to thermal images.
Unlike existing methods that rely on multimodal data with salient visual features, ThermoNeRF combines RGB and thermal features to learn consistent scene geometries across both sensor modalities, while avoiding contamination from each modality in the rendered views.
Concurrently to our work, other studies have investigated the use of NeRFs and thermal imaging [16, 17, 40]. However, they either rely on high-quality thermal images or focus on adverse conditions—such as low-light environments [18]—to enhance the visual quality of generated images, typically evaluated using RGB-specific metrics such as the PSNR, SSIM, and LPIPS. For instance, Xu et al. [18] only evaluate visible light image synthesis, neglecting the quality of the resulting thermal images. In contrast, our work prioritizes the accuracy of predicted temperatures in the synthesized novel views, which is crucial for practical, real-world applications such as thermal modeling of buildings.
2.3 Thermal 3D Reconstruction for Building Retrofit
Both 3D reconstruction and thermal imaging have been employed for building evaluation and retrofit. However, most state-of-the-art methods require either expensive and complex software [4], extensive data-collection and cleaning efforts [41], or a combination of both [42, 43].
Most similar to our work, Ham and Golparvar-Fard [44] uses photogrammetry to reconstruct four indoor scenes and detect condensation issues and thermal resistance of the wall. However, their approach shares the same limitations as the photogrammetry-based methods discussed in Section 2.1.
In conclusion, while both 3D reconstruction and thermal imaging have been employed for building evaluation and retrofit purposes, current state-of-the-art methods often present significant complexities and requirements. Our research tackle those issues by presenting a more practical and accessible approach to thermal modeling of buildings. ThermoNeRF combines the use of sparse data as input (i.e., flexibility in the data collection) and joint multi-modal reconstruction that does not rely on thermal to RGB projection.
@@ 4 ThermoNeRF
In this section, we present ThermoNeRF, a NeRF model capable of learning an implicit scene representation in both thermal and RGB, from a sparse set of RGB and thermal images. To synthesize realistic and accurate RGB and thermal novel views, ThermoNeRF retains geometric features captured by the RGB modality, while ensuring that temperature estimates are independent of scene color variations and details, and that they have a unique value for a given 3D point. Thus, our design choices are guided by two properties of thermal images: 1) thermal images are inherently soft and textureless, 2) surface temperatures are independent of the viewing direction.
Refer to Fig. 2 for a flowchart of the method; implementation of the model and evaluations can be found online333https://github.com/Schindler-EPFL-Lab/thermo-nerf.
@@ 4.1 Thermal Image Rendering
As seen in the thermal images of Fig. 3, the ghosting effect makes thermal images soft and textureless. Consequently, while textures and sharp geometric features in RGB images allow NeRF models to accurately estimate scene density, NeRF struggles to learn densities solely from thermal images—see Fig. 4(b) for experimental results of training Nerfacto on thermal images only. Therefore, a single density is shared by the modalities, enabling the network to construct a geometric representation informed by RGB information while ensuring geometric consistency with the thermal modality. This common density network is defined in the same way as in Section 3:
Since the aim of ThermoNeRF is to synthesize images that faithfully represent temperature at the surface, details present in the RGB images (such as edges, contrast, texture) should not “leak” into the synthetic thermal views, and colors in the RGB image should not influence the value of the synthesized temperatures in thermal images. Therefore, we propose to use separate MLPs to estimate each modality: predicts temperatures along each ray, while predicts RGB values.
Furthermore, as depicted in Fig. 3, color information is view-dependent—e.g., due to non-Lambertian effects—which is why, as seen in Eq. 1, takes as input , and to account for view direction. Thermal images are also view-dependent, not due to non-Lambertian effects but due to thermal reflections of nearby heat sources—for example, ceramic, metal, and glass are excellent thermal reflectors. However, in the context of non-destructive testing of infrastructure and infrastructure modeling, it is more interesting to reconstruct an image of the “real” temperature at the surface, rather than a realistic reconstruction of the scene influenced by thermal reflections, i.e. temperature should be invariant with respect to the viewing direction. Thus, only takes the intermediate features of as input, disregarding , and .
@@
where denotes the predicted temperature values and A is the appearance embedding.
@@ 4.2 Loss Functions
We use separate reconstruction loss functions for each modality. The final reconstruction loss is expressed as the sum of the Mean Square Error (MSE) losses of the RGB and thermal outputs, respectively denoted and . Similar to Nerfacto, we add the interlevel and distortion losses—respectively and —which were initially used in MIP-NeRF 360 [46], to optimize the proposal sampler and to reduce distortions respectively. Therefore, our final loss is defined as follows:
While we tried different values for and , we empirically found to be the best value and will thus use it in the rest of this paper.
@@ 5 ThermoScenes Dataset
*Due to the variations of atmospheric radiation in the sky, the readings of thermal cameras are subject to systematic inaccuracies [48] and the measurements are outside the camera’s operating range. We make available the raw data but, in our experiments, the temperature values under the minimum of the operating range of the camera (-20°C) are clipped to -20°C.
To evaluate RGB+thermal reconstruction and novel view synthesis methods, we have collected ThermoScenes, a new dataset comprising paired RGB and thermal images for 16 scenes (eight building facades and eight other everyday scenes) with diverse temperature ranges. A FLIR One Pro LT [49]—a dual RGB and thermal camera mounted on an iPhone and capable of simultaneously capturing aligned pairs of RGB and thermal images—was used to collect the data. The FLIR One Pro LT thermal camera operates within a range of -20 °C to 120 °C, with a thermal accuracy of 3 °C. In our study, evaluating the quality of 3D thermal reconstruction requires not only understanding the thermal accuracy but also the thermal precision of the camera. We assess the precision to be approximately C. To calculate the precision of our thermal camera, we fix the camera at a position and capture 20 consecutive images of a RaspberryPi connected to power for at least 10 minutes to ensure thermal stability. Our precision is then the average standard deviation of the temperatures of each pixel across the 20 images.
Furthermore, the FLIR software generates thermal images by overlaying thermal information onto RGB features—such as edges—using Multi-Spectral Dynamic Imaging (MSX) [50]. Instead, we provide and use raw thermal data extracted from MSX images using the FLIR Image Extractor library [51].
Table 1 provides a summary of the scenes comprising the ThermoScenes dataset, where each scene is illustrated by two example views in RGB and thermal respectively. The test set for each scene is uniformly sampled across the collected camera poses. The minimum and maximum measured temperatures are also reported. It should be noted that in the Building (spring), Dorm 1, MED, and INR building scenes, the sensor outputs values outside of its operating range due to the variations of atmospheric radiation in the sky, where the readings of thermal cameras are subject to systematic inaccuracies [48]. Given our focus on the object of interest—the building facade—we limit these values to the minimum of the operating range of the camera (-20°C) in our experiments. However, for other applications where more precise temperature estimates for the sky might be needed, we provide the raw data to allow for potential post-processing (see Kruczek [48] and Zhao et al. [52]).
ThermoScenes is publicly available444http://dx.doi.org/10.5281/zenodo.10835108 and we plan to gradually increase the size of the dataset. A tutorial on how to collect, process, and train data on ThermoNeRF has also been released.555https://github.com/Schindler-EPFL-Lab/thermo-nerf/blob/main/thermo_scenes/docs/Collect_new_dataset.md
| Metric | Method |
MED | Dorm 1 | Dorm 2 | INR | BI | Avg | ||||||
| PSNR | 20.30 | 22.80 | 23.88 | 19.26 | 18.65 | 19.42 | 18.8 | 22.95 | 20.77 | |||||||
| 20.60 | 28.47 | 27.74 | 20.58 | 32.44 | 26.64 | 21.05 | 27.15 | 26.24 | ||||||||
| ThermoNeRF | 26.63 | 28.75 | 33.79 | 22.04 | 34.10 | 29.94 | 22.45 | 28.83 | 28.62 | |||||||
| SSIM | 0.91 | 0.87 | 0.94 | 0.93 | 0.92 | 0.92 | 0.87 | 0.85 | 0.90 | |||||||
| 0.89 | 0.89 | 0.95 | 0.93 | 0.95 | 0.95 | 0.86 | 0.86 | 0.91 | ||||||||
| ThermoNeRF | 0.92 | 0.88 | 0.97 | 0.93 | 0.96 | 0.95 | 0.88 | 0.87 | 0.92 | |||||||
| MAE ∘C | 6.36 | 1.80 | 1.29 | 4.95 | 4.37 | 3.93 | 1.81 | 1.70 | 3.09 | |||||||
| 6.54 | 0.86 | 1.00 | 3.62 | 0.63 | 0.77 | 1.31 | 1.04 | 1.79 | ||||||||
| ThermoNeRF | 1.88 | 0.66 | 0.31 | 3.10 | 0.38 | 0.75 | 1.28 | 0.74 | 1.04 | |||||||
| MAE ∘C | 6.74 | 1.87 | 1.56 | 4.62 | 4.06 | 3.65 | 2.16 | 1.81 | 3.11 | |||||||
| 6.59 | 0.89 | 0.97 | 3.80 | 0.68 | 3.65 | 1.63 | 1.00 | 2.17 | ||||||||
| ThermoNeRF | 2.40 | 0.76 | 0.35 | 3.38 | 0.46 | 0.77 | 0.93 | 0.76 | 1.13 |